### v2

# Giant and robust Josephson diode effect in multiband topological nanowires

Abstract
[…]
Introduction.– The Majorana bound states (MBSs) have attracted significant interest for obeying non-Abelian statistics, and can serve as a basic building block for fault tolerant quantum computing Alicea (2012); Kitaev (2003); Nayak et al. (2008); Ivanov (2001); He et al. (2020); Chan et al. (2017). A leading platform for realizing MBSs is the proximitized nanowire Lutchyn et al. (2010); Oreg et al. (2010); Sau et al. (2010); Laubscher and Klinovaja (2021); Stanescu et al. (2011); Lutchyn et al. (2011); Qiao et al. (2024), where a semiconductor wire with strong spin-orbit interaction (SOI) is coupled to a superconductor under a magnetic field. This setup can host MBSs at two ends of the nanowire, with experimental signatures such as zero-bias conductance peaks Prada et al. (2020); Sasaki et al. (2011); Deng et al. (2012); Das et al. (2012); Mourik et al. (2012); Churchill et al. (2013) and the fractional $4\pi$-periodic Josephson effect Kwon et al. (2004a); Fu and Kane (2009); Kwon et al. (2004b); Laroche et al. (2019); Wiedenmann et al. (2016). The latter provides a key tool for probing MBSs, reflecting their unique phase behavior in Josephson junctions. Despite significant progress Wang et al. (2018); Albrecht et al. (2016); Deng et al. (2016); Quantum et al. (2025); Ren et al. (2019), obtaining unambiguous confirmation of MBSs remains a persistent challenge that continues to drive ongoing research.
Diodes, which allow current to flow mainly in one direction, are essential in modern electronics and have played a key role in the development of the information age. While conventional diodes suffer from Joule heating due to resistance, superconducting diodes provide a dissipation free alternative. This advantage has spurred rapid interest in the superconducting diode effect (SDE). The SDE is generally divided into intrinsic Ando et al. (2020); Miyasaka et al. (2021); Yuan and Fu (2021); He et al. (2022); Narita et al. (2022); Yuan and Fu (2022a); Daido et al. (2022); Zinkl et al. (2022); Hou et al. (2023); Anh et al. (2024); Mao et al. (2024a) and Josephson diode effects (JDE) Hu et al. (2007a); Misaki and Nagaosa (2021a); Wang et al. (2022); Davydova et al. (2022a); Wu et al. (2022a); Baumgartner et al. (2022a); Pal et al. (2022a); Zhang et al. (2022), with the latter drawing particular attention. When both inversion and time-reversal symmetries are broken Liu and Lobos (2013); Liu (2012), the forward ($I_{c}^{+}$) and reverse ($I_{c}^{-}$) directional critical currents are generally unequal Liu and Lobos (2013), a key mechanism underlying the SDE. In JDE, the Andreev bound states (ABSs) in the junction region contribute to the supercurrent Sauls (2018); Mizushima and Machida (2018); Hays et al. (2018); Tosi et al. (2019); Nichele et al. (2020). Recent experiments have observed the JDE in various systems Ciaccia et al. (2023); Banerjee et al. (2023); Matsuo et al. (2023); Trahms et al. (2023); Pal et al. (2022b); Baumgartner et al. (2022b); Díez-Mérida et al. (2023); Wu et al. (2022b), sparking significant interest. Alongside these experimental advancements, theoretical frameworks have been developed for the JDE behavior Cheng and Sun (2023); Davydova et al. (2022b); Hu et al. (2007b); Misaki and Nagaosa (2021b); Lu et al. (2023); Souto et al. (2022); Halterman et al. (2022); Kokkeler et al. (2022); Fominov and Mikhailov (2022); Kopasov et al. (2021); Gupta et al. (2023); Sun et al. (2023); Jiang et al. (2022); Tanaka et al. (2022); Maiani et al. (2023); Hu et al. (2023); Misaki and Nagaosa (2021b); Steiner et al. (2023); Souto et al. (2022); Pillet et al. (2023); Hess et al. (2023); Costa et al. (2025); Fukaya et al. (2024); Debnath and Dutta (2025), of particular interest is the JDE realized in topological nanowires hosting MBSs. Previous studies have focused on ideal one-dimensional (1D) models Liu et al. (2024); Legg et al. (2023); Mondal et al. (2025), whereas realistic systems with multiband Stanescu et al. (2011); Lutchyn et al. (2011); Liu et al. (2012) have been largely overlooked. Such systems may exhibit distinct mechanisms that can significantly modify the diode efficiency.
[…]
In this Letter, we predict a giant and robust JDE, together with a novel mechanism uncovered here, for Majorana nanowires in the multiband regime. The multiband topological nanowire hosts both MBSs and ABSs, which we show produce a large JDE by balancing the contributions from the fractional and conventional Josephson effects, with the diode efficiency remaining substantial in the deep topological phase. Importantly, we uncover a novel mechanism that the spin-parity band exchange among subbands leads to a robust plateau of high diode efficiency as tuning the external magnetic field. Our findings emphasize the critical role of the spin-parity subband engineering in optimizing the JDE, and provide a realistic method to identify topological phases for MBSs.
The multiband model.–We start with the multiband model for the quasi-1D nanowire placed on the $s$-wave superconductor to generate the proximity-induced superconductivity, forming a Josephson junction as depicted in Fig. 1(a). The system has finite dimensions $L_{x}$, $L_{y}$, and $L_{z}$ along the three spatial directions, which can be effectively modeled using square or harmonic trapping potentials. In typical nanowire material, the system is strongly confined along the $y$ and $z$ directions. To facilitate the theoretical description, we consider the rectangular cross section with a width of $L_{y}\sim 10^{2}\,\text{nm}$ and a thickness of $L_{z}\sim 10^{0}\,\text{nm}$), resulting in a reasonable approximation of $L_{z}\ll L_{y}\ll L_{x}$. Consequently, we limit our analysis to the lowest eigenstates in the $z$-direction, setting $N_{z}=1$ Stanescu et al. (2011); Lutchyn et al. (2011). As for the $y$-direction, we focus on cases that a few transverse eigenmodes in this direction can be occupied (e.g. with the number of orbits $N_{y}\leq 3$, see Fig. 1(b)). The resulting effective lattice Hamiltonian takes the form $H_{\mathrm{nw}}=H_{0}+H_{\text{SOI}}+H_{B}+H_{\Delta}$, where $H_{\mathrm{SOI}}$ denotes the Rashba SOI, $H_{B}$ accounts for the Zeeman splitting due to an external magnetic field, and $H_{\Delta}$ represents the induced pairing term. The explicit form of each contribution is given by
$H_{0}=-t_{x}\sum_{\boldsymbol{i}}(c_{\boldsymbol{i}+\boldsymbol{\delta}_{x}}^{\dagger}c_{\boldsymbol{i}}+\text{h.c.})-\sum_{\boldsymbol{i}}(\mu_{\boldsymbol{i}}-2t_{x})c_{\boldsymbol{i}}^{\dagger}c_{\boldsymbol{i}},$
$H_{\text{SOI}}=\mathrm{i}\sum_{\boldsymbol{i}}\left[c_{\boldsymbol{i}+\boldsymbol{\delta}_{x}}^{\dagger}(\alpha_{x}\hat{\sigma}_{y})c_{\boldsymbol{i}}-c_{\boldsymbol{i}+\boldsymbol{\delta}_{y}}^{\dagger}(\alpha_{y}\hat{\sigma}_{x})c_{\boldsymbol{i}}\right]+\text{h.c.},$
$H_{B}=\sum_{\boldsymbol{i}}c_{\boldsymbol{i}}^{\dagger}(h_{x}\hat{\sigma}_{x}+h_{y}\hat{\sigma}_{y})c_{\boldsymbol{i}},$
$H_{\Delta}=\sum_{\boldsymbol{i}}\left(\Delta_{\boldsymbol{i}}c_{\boldsymbol{i}\uparrow}^{\dagger}c_{\boldsymbol{i}\downarrow}^{\dagger}+\text{h.c.}\right),$ (1)
with electron creation operator denoted by the spinor $c_{\boldsymbol{i}}^{\dagger}=(c_{\boldsymbol{i}\uparrow}^{\dagger},c_{\boldsymbol{i}\downarrow}^{\dagger})$ and $\boldsymbol{i}=(i_{x},i_{y})$. Note that the index $i_{x}$ denotes the real site position along the $x$-direction, while $i_{y}=1,2,...,N_{y}$ is an effective position denoting transverse modes in the $y$ direction and characterizes an orbital degree of freedom. The Hamiltonian consists of four contributions: The free-particle term $H_{0}$ describes nearest-neighbor hopping $t_{x}$ along $\boldsymbol{\delta}_{x}=(1,0)$ and includes an $i_{y}$-site-dependent (actually transverse-mode dependent) chemical potential $\mu_{\boldsymbol{i}}=\mu_{0}-i_{y}^{2}(\pi\hbar)^{2}/(2m^{*}L_{y}^{2})$ Further details are provided in the Supplementary Material , where the energy of transverse modes is approximated by that in a 1D infinite square well Lutchyn et al. (2011), and $m^{*}$ is the effective mass of the electron. The SOI term $H_{\text{SOI}}$ incorporates the longtitutional SOI with intensity $\alpha_{x}$ and the effective transverse SOI between different orbits with intensity $\alpha_{y}\propto L_{y}^{-1}$ , which is nonzero only when $\boldsymbol{\delta}_{y}$=(0,1) Further details are provided in the Supplementary Material , with the Pauli matrices $\hat{\sigma}_{i}(i=x,y,z)$ on spin space. The term $H_{B}$ describes a uniform magnetic field $\boldsymbol{h}=(h_{x},h_{y})$ that applied throughout the entire nanowire, where $h_{x}$ opens a Zeeman gap in the spectrum, while $h_{y}$ shifts the Fermi surface along the $x$ direction, breaking inversion symmetry Liu and Lobos (2013); Yuan and Fu (2022b). Finally, $s$-wave pairing is depicted by $H_{\Delta}$, with $\Delta_{\boldsymbol{i}}$ nonzero only in the superconducting leads, carrying a phase difference $\phi$ across the junction: $\Delta_{\mathrm{R}}=\Delta_{\mathrm{L}}e^{i\phi}=\Delta e^{i\phi}$.
Balance between fractional and conventional Josephson currents.– In the multiband regime, the Fermi level crosses multiple Fermi points, such that the topological nanowire may generically host both MBSs and ABSs in the boundary [Fig. 1(a)], which in the Josephson junction give rise fractional ($4\pi$-periodic) and conventional ($2\pi$-periodic) Josephson current components, denoted $I_{4\pi}$ and $I_{2\pi}$ [Fig. 1(d)], respectively. The total supercurrent at zero temperature versus superconducting phase difference $\phi$ reads Beenakker (1991)
$I(\phi)=\frac{2e}{\hbar}\frac{\partial E}{\partial\phi},\,E=-\frac{1}{2}\sum_{E_{i}\geq 0}E_{i}(\phi).$ (2)
where the sum runs over all positive-energy eigenstates. Fig. 1(d) displays the total current, $I_{\text{tot}}=I_{4\pi}+I_{2\pi}$ (red line), which develops a global maximum distinct from its minimum. Consequently, the critical currents become asymmetric, $I_{c}^{+}\neq I_{c}^{-}$, with $I_{c}^{\pm}=\max[\pm I_{\text{tot}}(\phi)]$. This asymmetry constitutes direct evidence of the Josephson diode effect, characterized by the efficiency $\eta=(I_{c}^{+}-I_{c}^{-})/(I_{c}^{+}+I_{c}^{-})$. In this multiband regime, the JDE originates from the competition between the fractional current $I_{4\pi}$ and the conventional component $I_{2\pi}$.
We refer to realistic parameters from experiments on InAs and InSb nanowires placed on Nb or Al superconductors Lutchyn et al. (2011). The nanowire width is $L_{y}=130\,\text{nm}$ and the effective lattice constant in the $x$-direction $a_{x}$ is taken to be $4\,\text{nm}$. Then $m^{*}=0.04m_{e}$, with SOI strength $\alpha_{x}=0.1\,\text{eV}\cdot\mathring$$\text${A}} and BCS pairing potential $\Delta=2.4\,\text{K}$ ($0.2\,\text{meV}$). The corresponding tight-binding parameters for numerical study are $t_{x}=200\Delta$, $\alpha_{x}=6\Delta$, $\alpha_{y}=1.3\Delta$, and let $\Delta=1$ for convenience. Further, to enable a deep topological phase regime for our study, the Fermi surface can be adjusted to be halfway between the two eigenmodes at $k_{x}=0$, which determines the value of $\mu_{0}$. We consider a short junction limit with a normal region length $N_{\text{N}}=2$, and the Josephson current is computed using the recursive Green’s function method Song et al. (2016); Mao et al. (2024b):
$I=-e\left\langle\frac{\mathrm{d}\hat{N}_{\text{L}}}{\mathrm{d}t}\right\rangle=\frac{e}{\hbar}\mathrm{Tr}\left\{\Gamma_{z}\left[{G}_{\text{10}}^{<}(t,t)\hat{T}-\hat{T}^{\dagger}{G}_{\text{01}}^{<}(t,t)\right]\right\}.$ (3)
Here, the particle number on the left lead is defined as $\hat{N}_{\text{L}}=\sum_{i_{x}\leq 0,i_{y}}\sum_{\sigma=\uparrow,\downarrow}c_{\boldsymbol{i}\sigma}^{\dagger}c_{\boldsymbol{i}\sigma}$, with $\Gamma_{z}=\sigma_{z}\otimes I_{N_{y}\times N_{y}}$ and $\hat{T}$ is the hopping matrix connecting the left superconductor to the normal region, $G_{\text{NL}}^{<}(t,t)$ is the lesser Green’s function. For simplicity, we often set $e=\hbar=1$.
Figure 2: Relatively high and stable efficiency $\eta$ of the system populating 3 spinful subbands in the region of strong $h_{x}$ under various parameters. (a) $\eta$ vs $h_{x}$ for typical $\alpha_{x}$’s when 1 subband (calculated with $N_{y}=1$, ideal 1D model) or 3 subbands (evaluated with $N_{y}=2$) are occupied. (b) $\eta$ vs $h_{x}$ for realistic $L_{y}$’s. (c) $\eta$ vs $h_{y}$ for different $h_{x}$’s. (d) $\eta$ vs $\mu_{0}$ when other parameters are fixed. “Tri”, “Topo”, “Gapl” and “Vac” stand for “Trivial”, “Topological”, “Gapless” and “Vacuum”, respectively. Parameters: $\alpha_{x}$ is set to be 30 in panel (b)-(c), and $h_{y}=0.8$ in panel (a)-(c). In panel (d), $\alpha_{x}=10,\,h_{x}=5,\,h_{y}=0.7$.
Fig. 2 shows the superconducting diode efficiency versus the magnetic field $h_{x}$ for different system parameters like the SOI strength $\alpha_{x}$ (Fig. 2(a)), the nanowire’s transverse length $L_{y}$ (Fig. 2(b)), and $h_{y}$ (Fig. 2(c)). For the single band regime (see the colored dashed lines), we can find that the diode efficiency $\eta$ consistently exhibits a pronounced peak near the critical topological phase transition boundary $h_{x}^{c}=\sqrt{\Delta^{2}-h_{y}^{2}}$ Liu et al. (2024) and diminishes if tuning into deep topological and trivial regimes. In the trivial (topological) regime, the JDE are primarily contributed from the higher harmonic terms of conventional (fractional) Josephson currents $I_{2\pi}$ ($I_{4\pi}$) due to the ABSs (MBSs) Legg et al. (2023), for which the diode effect is weak. Only near the phase transition boundary, which enables a competition between the two types of Josephson currents, an enhanced Josephson diode effect is resulted.
The multiband regime shows a novel new feature that the high diode efficiency is generically obtained in the deep topological phase regime. Within the multiband topological phase, the contributions from ABSs and MBSs coexist, leading to a persistent competition between $I_{4\pi}$ and $I_{2\pi}$ across the entire topological regime. Consequently, the regime of high diode efficiency extends over most of the topological phase region [Fig. 2(a)-(c)]. Moreover, the maximum position of diode efficiency varies depending on system parameters, but is always obtained from the existing balance between $I_{4\pi}$ and $I_{2\pi}$ in deep topological phase. In Fig. 2(d), increasing the chemical potential drives the Fermi level across one, two, and three subbands, defining regimes I, II, and III. Region $\mathrm{I}$ features a low efficiency platform contributed from MBSs and region $\mathrm{II}$ exhibits weak diode effect dominated by ABSs. In contrast, in region $\mathrm{III}$ a high diode efficiency is driven by the coexistence of MBSs and ABSs.
Figure 3: High diode efficiency plateau by spin-parity subband exchanges mechanism. (a) Energy bands before, during, and after the spin-parity band exchange in the 2-orbit ($N_{y}=2$) model with 3 occupied subbands. Colored bands denote different spin-parity subbands, and colored arrows mark the Fermi point shifts induced by the inversion-symmetry-breaking field $h_{y}$. (b) Diode efficiency $\eta$ as a function of $h_{x}$, where the background colors correspond to the band structures shown in (a). (c) Energy bands showing the spin-parity band exchange in the three-orbital ($N_{y}=3$) model with five occupied subbands. (d) $\eta$ versus $h_{x}$, with color segments matching the band structures in (c). Parameters: $\alpha_{x}=10,\,h_{y}=0.8$.
Spin-parity band exchange mechanism.–A more intriguing discovery is that we predict a robust high diode efficiency plateau with a novel spin-parity band exchange mechanism that is obtained only in the multiband regime. The spin-parity $\mathbb{P}_{s}$ of a certain subband is defined as the sign of the eigenvalue of the spin-coupled term of the Hamiltonian $h_{x}\hat{\sigma}_{x}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})\hat{\sigma}_{y}$ at $k_{y}=0$, where the interband SOI vanishes. As illustrated in Fig. 3, the bands with $\mathbb{P}_{s}=+1$ and $\mathbb{P}_{s}=-1$ are plotted in red and blue color, respectively. The relative positions of normal subbands (with $\Delta=0$) with different spin parities can be tuned by $h_{x}$ and shows a crucial impact on the diode effect. As $h_{x}$ increases, the subbands with opposite spin parities shift oppositely in energy, leading to the spin parity exchange of the two types of subbands when $h_{x}$ exceeds a threshold. We consider first the topological phase for $N_{y}=2$, with three normal subbands crossing Fermi energy [Fig. 3(a)]. We can see that at $h_{x}=0.7\Delta$ with $h_{y}=0.8\Delta$, on Fermi level there are two blue bands with $\mathbb{P}_{s}=-1$ and one red band with $\mathbb{P}_{s}=+1$. At $h_{x}\approx|\mu_{1}-\mu_{2}|/2\approx 3.5\Delta$, the upper two subbands begin to swap positions, and the full exchange is completed at $h_{x}\geq 4.7\Delta$. Then the spin parity configuration is robust for further increasing $h_{x}$. The diode efficiency in Fig. 3(b) highlights three distinct regions corresponding to the band exchange shown on the left. Prior to the band exchange, the diode efficiency remains to be low. At the band exchange region, a sharp enhancement in the diode efficiency emerges. After the complete exchange of the two subbands with different spin parities, a robust high diode efficiency plateau is obtained.
The predicted high diode efficiency plateau is not limited to $N_{y}=2$, but a generic result applicable to the multiband regime. To illustrate this, we consider further the case of $N_{y}=3$, in which the topological phase is obtained by setting the Fermi level to cross five subbands, including two red subbands and three blue subbands [Fig. 3(c)]. With increasing $h_{x}$, two spin parity exchange transitions are obtained, respectively at $h_{x}\approx 6.2\Delta$ and $h_{x}\approx 8.8\Delta$. Around the transition points, the diode efficiency exhibits local maximum or minimum values depending on details. In comparison, after the second transition, the spin parity configuration is stabilized, and the robust high diode efficiency plateau is generically obtained [Fig. 3(d)].
Figure 4: Underlying mechanism of high diode efficiency plateau after the spin-parity subband exchange. (a) Dependence of Fermi momenta $(k_{\textrm{F}})_{i_{y}}^{\pm}a_{x}$ on field strength $h_{x}$ for $N_{y}=2$ with three occupied subbands and the same parameters as Fig. 3(a). (b) Fermi momentum shifts $\Delta k_{\textrm{F}}=k_{\textrm{F}}(h_{y})-k_{\textrm{F}}(h_{y}=0)$. The yellow line denotes a dimensionless ratio $\xi$ that quantifies the competition between left- and right-shifted Fermi points. Parameters: $\alpha_{x}=10$, $h_{y}=0.8$.
The giant diode efficiency plateau implies a novel mechanism featured by the spin parity configuration of the subbands. As illustrated in Fig. 3(a), the Fermi level intersects three subbands. For $h_{y}=0$, the bands remain symmetric about $k_{x}=0$. With $h_{y}\neq 0$, the 1D dispersions take the form $E_{i_{y}}^{\pm}(k_{x})=2t_{x}(1-\cos k_{x}a_{x})-\mu_{i_{y}}\pm\sqrt{h_{x}^{2}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})^{2}},$ with the orbit and spin-parity being labeled by $i_{y}$ and $\pm$, respectively. Therefore, upon applying $h_{y}$, the net shift of the two Fermi points in a given subband is
$(\Delta k_{\textrm{F}})_{i_{y}}^{\pm}\approx\mp\frac{2}{\hbar|v_{\textrm{F}}|}{\left[1+\left(\frac{h_{x}}{2\alpha_{x}\sin k_{\textrm{F}}a_{x}}\right)^{2}\right]}^{-\frac{1}{2}}h_{y}.$ (4)
This expression quantifies the directional deformation of Fermi points, where the red bands ($\mathbb{P}_{s}=+1$) shift leftward and the blue bands ($\mathbb{P}_{s}=-1$) shift rightward [Fig. 3(a)]. Fig. 4(a,b) shows the Fermi point positions and their shifts for the three occupied bands with $N_{y}=2$ as a function of the magnetic field $h_{x}$. We find that $(\Delta k_{\textrm{F}})_{1}^{+}$ and $(\Delta k_{\textrm{F}})_{2}^{-}$ initially decrease, cross zero, and eventually saturate. The zero crossing signifies an exchange of the Fermi points. These Fermi point shifts are induced by the transverse field $h_{y}$. A key feature is that the dimensionless ratio $\xi=2|(\Delta k_{F})_{1}^{+}|/[(\Delta k_{F})_{1}^{-}+(\Delta k_{F})_{2}^{-}]$, which quantifies the relative shift between left- and right-shifted total Fermi momenta, saturates to a stable plateau after the spin-parity exchange [the yellow line in Fig. 4(b)]. This saturation corresponds to the emergence of high diode efficiency plateau. In this regime, the two blue bands with spin-parity $\mathbb{P}_{s}=-1$ exhibit nearly identical and constant Fermi point displacements, which mainly contribute the conventional Josephson current $I_{2\pi}$. In contrast, the remaining red subband predominantly contributes to the fractional Josephson current $I_{4\pi}$ related to MBSs. The saturation of $\xi$ signals the balance between the two types of Josephson currents and the high diode efficiency plateau.
We also note that the balance is resulted from the competition between the Josephson current originating from the higher-curvature red band and that originating from the lower-curvature blue subbands. In general, through the spin-parity band exchange, we can combine the $N_{y}$ lower-curvature bands of identical spin parity to balance the $N_{y}-1$ opposite spin-parity bands with higher curvature. Such a mechanism provides a generic practical route to optimize the Josephson diode effect with high and robust efficiency in the multiband topological regime.
Conclusion.–We have shown that multiband Majorana nanowires provide a realistic and robust platform for realizing the giant Josephson diode effect. The coexistence of MBS and ABS in the multiband regime leads to a persistent competition between the fractional and conventional Josephson currents, sustaining large diode efficiency deep into the topological phase. Especially, we uncovered a novel spin-parity band exchange mechanism unique to the multiband regime, which realizes a robust high efficiency plateau. These results demonstrate subband engineering as a practical route to optimize superconducting diode effect and offer a new probe for identifying topological phase in Majorana nanowire systems.
Acknowledgments.– The authors are thankful to Yue Mao and Yu-Hao Wan for fruitful discussions. This work was supported by National Key Research and Development Program of China (2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12425401, No. 12261160368, and No. 11921005), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), and Shanghai Municipal Science and Technology Major Project (No. 2019SHZDZX01).
Supplementary Material: Giant and robust Josephson diode effect in multiband topological nanowires
S-1 From quasi-1D nanowire to multiband Hamiltonian
In this section, we derive the effective tight-banding multiband Hamiltonian from a realistic quasi-1D nanowire, and find proper values for parameters for numerical calculation. Specifically, we consider a semiconductor quantum well based on an InAs-Al/Sb heterostructure. As reported in Refs. Stanescu et al. (2011); Lutchyn et al. (2011), it exhibits an effective electron mass of $m^{*}=0.04\,m_{e}$, and a spin-orbit interaction (SOI) strength of approximately $(\alpha_{x})_{\text{C}}=0.1\,\text{eV}\cdot\mathring$$\text${A}} (for the continuous model). Typical system dimensions are $L_{z}<10\,\text{nm}$, $L_{y}\approx 130\,\text{nm}$, and $L_{x}$ that extend over several microns, validating the approximation $L_{z}\ll L_{y}\ll L_{x}$.
We first derive the hopping term $t_{x}$ and the spin-orbit interaction term $\alpha_{x}$ along the $x$-direction, which correspond to the parameters of an ideal one-dimensional nanowire. The horizontal lattice constant is set to be $a_{x}=4\,\text{nm}$. For a simple lattice system with only spin-independent hopping and Rashba SOI, it is crucial to ensure that both the continuous model (described by $m^{*}$ and $(\alpha_{x})_{\text{C}}$) and the tight-banding model (with the nearest-neighbor hopping element $t_{x}$ and $(\alpha_{x})_{\text{TB}}$) accurately capture the low-energy physics near the band bottom. To establish a clear correspondence between these descriptions, we expand the electronic dispersion around $k_{x}=0$. By retaining terms up to the first-order in $k_{x}$, we extract the SOI parameter for the tight-binding model, while the hopping term is determined from the second-order expansion:
$(\alpha_{x})_{\text{C}}k_{x}\approx 2(\alpha_{x})_{\text{TB}}\sin(k_{x}a_{x}),$ $\,\quad\alpha_{x}=(\alpha_{x})_{\text{TB}}=\frac{(\alpha_{x})_{\text{C}}}{2a_{x}}=1.2\,\text{meV},$
$\frac{\hbar^{2}k_{x}^{2}}{2m^{*}}\approx 2t_{x}-2t_{x}\cos(k_{x}a_{x}),$ $\,\quad t_{x}=\frac{\hbar^{2}}{2m^{*}a_{x}^{2}}=40\,\text{meV}.$ (S1)
Next, we estimate the parameters $\mu_{\boldsymbol{i}}=\mu_{i_{y}}$ and $\alpha_{y}$ to qualitatively describe multiple orbits. The variation in the chemical potential across different orbits arises from the confinement along the $y$-direction. For simplicity, we model the system in the $y$-direction as an infinite potential well, leading to the following energy spectrum and corresponding eigenstates:
$E(i_{y})=i_{y}^{2}\cdot\frac{\pi^{2}\hbar^{2}}{2m^{*}L_{y}^{2}},\,\quad\psi_{i_{y}}(y)=\sqrt{\frac{2}{L_{y}}}\sin\frac{i_{y}\pi y}{L_{y}}.$ (S2)
Here, $i_{y}$ is a positive integer that labels the energy levels. These energy levels determine the variation in the chemical potential among the subbands, while the background chemical potential $\mu$ can be uniformly tuned via a metallic gate in experiments:
$\mu_{i_{y}}=-\mu_{0}+E(i_{y})=-\mu_{0}+i_{y}^{2}\cdot(\frac{\pi a_{x}}{L_{y}})^{2}t_{x}.$ (S3)
From the above, we can define $V=E(1)=\pi^{2}(a_{x}/L_{y})^{2}t_{x}=0.5\,\text{meV}$ to reflect the characteristic energy difference between orbits in the $y$ dimension in the effective lattice model, with higher energy levels following $E(i_{y})=i_{y}^{2}\cdot E(n_{y}=1)=i_{y}^{2}V$. Meanwhile, the SOI in the $y$-direction can be incorporated by the second quantization:
$2\alpha_{x}a_{x}\braket{\psi_{i_{y}}|\partial_{y}|\psi_{i_{y}^{\prime}}}=2\alpha_{x}a_{x}\int_{0}^{L_{y}}{\frac{2}{L_{y}}}\sin\frac{i_{y}\pi y}{L_{y}}\partial_{y}\sin\frac{i_{y}^{\prime}\pi y}{L_{y}}\mathrm{d}y=\frac{2\alpha_{x}a_{x}}{L_{y}}A_{i_{y}i_{y}^{\prime}}=\alpha_{y}\frac{A_{i_{y}i_{y}^{\prime}}}{A_{21}}.$ (S4)
One can calculate $\alpha_{y}=0.2\,\text{meV}$, and $A_{i_{y}i_{y}}=0$, $A_{21}=8/3$, $A_{31}=0$, $A_{32}=24/5$.
To simplify numerical calculations, physical quantities with energy dimensions are rescaled to dimensionless units. Given the significant contribution of Majorana bound states to the supercurrent, it is convenient to set the BCS pairing strength $\Delta=2.4\,\text{K}=0.2\,\text{meV}$ as unity in computations. Consequently, the parameters are chosen as $t_{x}=200$, $\alpha_{x}=6$, $V=E(n_{y}=1)=2.5$, and $\alpha_{y}=1.3$, ensuring consistency with the precision of $m^{*}$ and $(\alpha_{x})_{\text{C}}$. Considering possible candidate materials for the nanowire, one can adopt the SOI intensity of the same magnitude.
For the possible values of $h_{x}$ and $h_{y}$, we note that the effective Land¨¦ factor in an InSb-based semiconducting nanowire is $g_{\text{eff}}\approx 50$, with $t_{x}$ and $\alpha_{x}$ comparable to those in InAs. By adopting a dimensionless $h$ and normalizing $\Delta$ to unity in the computation, the corresponding magnetic field in SI units is given by $B=\Delta/(g_{\text{eff}}\mu_{\text{B}})\cdot h=0.07h\cdot\text{T}$. Thus, achieving the high-efficiency regime requires a magnetic field just below $1\text{T}$, which is experimentally available.
S-2 The recursive Green’s function
Figure S1: The tight binding model for the multiband nanowire. Here $N_{y}$ denotes the number of orbital channels included in the system. Each orbital is treated as a chain along the $x$-direction. The coordinate origin in the $x$-direction is defined at the interface between the left lead and the normal region.
We present the details of the recursive Green’s function method Mao et al. (2024a) employed to compute the supercurrent in the main text. The continuum system is discretized into a multichain tight-binding model, as schematically illustrated in Fig. S1. The system is partitioned into on-site terms, $H_{0,\text{L}/\text{R}/\text{N}}$, and coupling terms, $H_{x}$ and $H_{y}$. The on-site Hamiltonian consists of the left superconducting lead ($H_{0,\text{L}}$), the central normal region ($H_{0,\text{N}}$), and the right superconducting lead ($H_{0,\text{R}}$), each comprising $N_{y}$ tight-binding chains, where $N_{y}$ corresponds to the number of orbital channels considered. The chains are coupled via spin-orbit interactions described by $H_{y}$, while adjacent sites along the $x$-direction are connected through $H_{x}$. The explicit forms of these terms are given below.
$H_{0,\text{L}}$ $=-\mu_{s}\sigma_{0}\otimes\tau_{z}+h_{x}\sigma_{x}\otimes\tau_{z}+h_{y}\sigma_{y}\otimes\tau_{0}+\Delta\sigma_{y}\otimes\tau_{y}$
$H_{0,\text{N}}$ $=-\mu_{n}\sigma_{0}\otimes\tau_{z}+h_{x}\sigma_{x}\otimes\tau_{z}+h_{y}\sigma_{y}\otimes\tau_{0}$
$H_{0,\text{R}}$ $=-\mu_{s}\sigma_{0}\otimes\tau_{z}+h_{x}\sigma_{x}\otimes\tau_{z}+h_{y}\sigma_{y}\otimes\tau_{0}+\Delta\left(\cos\phi\cdot\sigma_{y}\otimes\tau_{y}+\sin\phi\cdot\sigma_{y}\otimes\tau_{x}\right)$
$H_{x}$ $=-t_{x}\sigma_{0}\otimes\tau_{z}+\text{i}\alpha_{x}\sigma_{y}\otimes\tau_{z}$ (S5)
$H_{y}$ $=\alpha_{y}\sigma_{x}\otimes\tau_{0}$
Here, $\sigma$ denotes the spin degree of freedom, while $\tau$ labels the particle-hole space. The three regions differ only in the superconducting term $\Delta$: it vanishes in the central normal region ($\Delta=0$) but is finite in the left and right topological regions. A superconducting phase difference $\phi$ exists between the left and right superconductors. All other parameters, including the hopping term, magnetic field, and spin-orbit coupling, remain identical across the three regions, as they are part of the same nanowire. Consequently, the spin-orbit coupling is uniform along both the $x$- and $y$-directions.
In Fig. S1, the system extends significantly in the $x$-direction, while its size in the $y$-direction is restricted to $N_{y}$ sites. For simplicity, each column is treated as a single unit, with the Hamiltonian of an individual column denoted as $H_{00,\alpha}$, where $\alpha=(\text{L},\text{N},\text{R})$ represents the left superconducting region, the normal region, and the right superconducting region, respectively. The coupling between neighboring columns in the $x$-direction is described by $H_{01,\alpha}$ with $\alpha=(\text{L},\text{N},\text{R})$. Notably, $H_{01,\alpha}$ is identical across these regions, as the coupling in the $x$-direction involves only hopping and spin-orbit coupling, which remain uniform throughout the nanowire. Then the Hamiltonians could be rewritten as
$H_{00,\alpha}=\begin{bmatrix}H_{0,\alpha}&H_{y}&&&&\\ H^{*}_{y}&H_{0,\alpha}&H_{y}&&&\\ &H^{*}_{y}&H_{0,\alpha}&H_{y}&&\\ &&&\ddots&&\\ &&&H^{*}_{y}&H_{0,\alpha}&H_{y}\\ &&&&H^{*}_{y}&H_{0,\alpha}\end{bmatrix},\qquad H_{01,\alpha}=\begin{bmatrix}H_{x}&0&&&&\\ 0&H_{x}&0&&&\\ &0&H_{x}&0&&\\ &&&\ddots&&\\ &&&0&H_{x}&0\\ &&&&0&H_{x}\end{bmatrix}$ (S6)
The current can be derived from the Green’s function
$I$ $=-e\left\langle\frac{\mathrm{d}\hat{N}_{\text{L}}}{\mathrm{d}t}\right\rangle=\frac{\text{i}e}{\hbar}\left\langle\left[\sum_{i_{x}<0,\sigma}\hat{c}_{i_{x},\sigma}^{\dagger}\hat{c}_{i_{x},\sigma},H\right]\right\rangle$ (S7)
$=\frac{\text{i}e}{\hbar}\sum_{\sigma}(t_{0}\langle\psi^{\dagger}_{0,\sigma}\psi_{1,\sigma}\rangle-t_{0}\langle\psi^{\dagger}_{1,\sigma}\psi_{0,\sigma}\rangle)$
$=\frac{e}{\hbar}\mathrm{Tr}\left\{\Gamma_{z}\left[{G}_{\text{10}}^{<}(t,t)\hat{T}-\hat{T}^{\dagger}{G}_{\text{01}}^{<}(t,t)\right]\right\}.$
We adopt $\Gamma_{z}=\text{diag}(1,1,-1,-1)$ and transform the lesser Green’s function into energy space according to ${G}_{NL}^{<}=\int\mathrm{d}\epsilon{G}^{<}(\epsilon)/2\pi$, where $\epsilon$ denotes the energy. In the energy representation, the lesser Green’s function is given by ${G}^{<}(\epsilon)=-f(\epsilon)[{G}^{r}(\epsilon)-{G}^{a}(\epsilon)]$, with $f(\epsilon)$ being the Fermi distribution function. The retarded Green’s functions, $G_{\text{01}}^{r},G_{\text{10}}^{r}$, are obtained via the Dyson equations $G_{\text{01}}^{r}=g_{\text{00}}^{r}\Sigma_{\text{01}}^{r}G_{\text{11}}^{r},G_{\text{10}}^{r}=G_{\text{11}}^{r}\Sigma_{\text{10}}^{r}g_{\text{00}}^{r}$, with $\Sigma_{\text{01}}^{r}=\hat{T},\Sigma_{\text{10}}^{r}=\hat{T}^{\dagger}$. Here, $g_{\text{00}}^{r}$ is the surface Green’s function of the left lead. ${G}_{\text{11}}^{r}$ is the Green’s function of the leftmost layer $(i_{x}=1)$ of the normal region, calculated by a recursive algorithm Mao et al. (2024a); Song et al. (2016). The advanced Green’s function satisfies ${G}_{\text{10}}^{a}=({G}_{\text{01}}^{r})^{\dagger},{G}_{\text{01}}^{a}=({G}_{\text{10}}^{r})^{\dagger}$. Then we need to derive the surface Green’s function. Based on the Hamiltonian derived above, the Green function satisfies a recursive equation shown below,
$\begin{pmatrix}G_{n+2,0}\\ G_{n+1,0}\end{pmatrix}=\begin{pmatrix}H_{01}^{-1}(EI-H_{00})&H_{01}^{-1}H_{10}\\ I&0\end{pmatrix}\begin{pmatrix}G_{n+1,0}\\ G_{n,0}\end{pmatrix}=\hat{T}\begin{pmatrix}G_{n+1,0}\\ G_{n,0}\end{pmatrix}$ (S8)
we know that $G_{n,0}$ in the limit $n\to\infty$ must approach to zero. This means that the powers of the matrix on the right also approach zero. In other words, $G_{0}$ is composed of the eigenvectors of $\hat{T}$ corresponding to eigenvalues smaller than 1, which means
$\binom{G_{1,0}}{G_{0,0}}=\binom{S_{1}}{S_{2}}A,$ (S9)
where $\binom{S_{1}}{S_{2}}$ is the eigenvector matrix of $\hat{T}$ corresponding to eigenvalues smaller than 1. From this, we quickly obtain
$G_{00}=(EI-H_{00}-H_{01}S_{1}S_{2}^{-1})^{-1}.$ (S10)
Using the surface Green’s function obtained, we employ a recursive algorithm to determine the retarded Green’s function for the leftmost column of the normal region. The Green’s function $G_{i}$ for each column satisfies the recursive relation
$G_{i}=(EI-H_{00}-H_{10}G_{i-1}H_{01})^{-1}.$ (S11)
S-3 High-Efficiency Plateau with Different Other Parameters
Figure S2: The diode efficiency $\eta-h_{x}$ curves of the 2-orbit model with 3 subbands partially filled under different $t_{x}$’s (a) and for the diverse $N_{\text{N}}$’s (b). Parameters: $t_{x}=200,\,\alpha_{x}=30,\,V=2.5,\,h_{y}=0.8$.
In this section, we demonstrate that the high-$\eta$ plateau at large $h_{x}$ is observed across a wide range of $t_{x}$ and $N_{\text{N}}$ values, providing further evidence of its universality. As shown in Fig. S2(a), the efficiency curve remains almost unaffected by variations in $t_{x}$. This insensitivity can be attributed to the facts that the orientations of spins at low energy points (like $k_{x}=k_{\textrm{F}}\sim 2\alpha_{x}/(t_{x}a_{x})$ and $k_{x}=0$) are almost invariant under changes of $t_{x}$, and the bouncing mode(s) within the normal region of a relatively short Josephson junction is approximately irrelavant to the effective mass. As a result, both the MBSs and the ABSs are insensitive to altering of $t_{x}$ here. Additionally, Fig. S2(b) shows that this enhancement persists even when the junction length reaches $\sim 10a_{x}\sim 40\,\text{nm}$. The observed reduction in the efficiency plateau for larger $N_{\text{N}}$ can be explained by differing decay rates: while MBSs exhibit exponential decay, ABSs decay polynomially Sauls (2018), leading to a diminishing effect as $N_{\text{N}}$ increases.
S-4 Decline of the Efficiency Plateau under Inappropriate Subband Filling
As shown in Fig. S3(b), when a 3-orbit ($N_{y}=3$) model with 3 subbands populated (more realistic and reasonable than the model of $N_{y}=2$ model with 3 subbands occupied) is adopted, the left part of the efficiency curve $\eta-h_{x}$ is almost the same as that in Fig. 3(b). This is reasonable, since the subbands with $i_{y}=3$ are high-energy and far away from the Fermi surface, causing weak effect to the 2-orbit part, as illustrated by the band dispersions in Fig. S3(a). The shifts of Fermi points also make a proof, where Fig. 4(a) and the left part of Fig. S3(c) ($h_{x}<6$) share the same feature.
When the parallel component of external magnetic field $h_{x}$ continues to enlarge, the high efficiency plateau will turn into a slope downwards in the middle of the blue region. This decline happens when the top blue subband and the middle red one exchanges, and arises from the change of speed of Fermi points’ shifting.
As $h_{x}$ keeps increasing, there exists the second critical value, where a sudden drop instead of rise in $\eta$ occurs. At this time, the lowest red subband and the highest blue one pass through each other, while the Fermi points closest to $k_{x}=0$ suddenly encounter spin-parity flip. After that, the $\eta-h_{x}$ curve enters the relative final yellow region and performs as a relatively low platform, which arises from the imbalance of bound states: both ABSs and MBSs come from red subbands with the same shifting direction. This can be directly shown by the right part in Fig. S3(c) ($h_{x}>10$).
Figure S3: Cases of 3-orbit model with 3 subbands occupied. (a) The dispersion relations under different $h_{x}$’s. (b) The diode efficiency $\eta$ versus $h_{x}$. (c)-(d) The curves of the Fermi wavevectors and the shifts in the Fermi points as functions of $h_{x}$. Parameters: $t_{x}=200,\,\alpha_{x}=10,\,V=2.5,\,\alpha_{y}=1.3,\,h_{y}=0.8$.
In conclusion, this is an example of inappropriate subband filling under a relatively large magnetic field. This provides a counter-argument to the conjecture we made in the text, that it may be better to populate $2N_{y}-1$ spinful subbands when $N_{y}$-orbit model is necessary and enough under a certain $h_{x}$. If a specific number of subbands are occupied but the $h_{x}$ is constantly increased, the high-efficiency platform will not extend infinitely.
S-5 Fermi Point Shifting for Model with More Orbits
Figure S4: (a) Fermi vectors with and without $h_{y}$ for $N_{y}=3$ with 5 subbands populated. (b) Shifts of Fermi Points. Parameters: $t_{x}=200,\,\alpha_{x}=10,\,V=2.5,\,\alpha_{y}=1.3,\,h_{y}=0.8$.
In this section, we detail how the diode efficiency is controlled by adjusting the magnetic field $h_{x}$ for the case of $3$-orbital degree of freedom and $5$ subband occupied. As shown in Fig. S4(a), five pairs of Fermi points are plotted, and their positions vary as $h_{x}$ increases. This process reveals two distinct spin-parity subband exchange regions, since there is always $|k_{\textrm{F}}|^{-}_{2}>|k_{\textrm{F}}|^{+}_{1}$ here because of adequate SOI, avoiding one exchange. Correspondingly, in Fig. S4(b), the Fermi point shift $\Delta k_{x}a_{x}$ crosses zero twice, indicating band exchanges happens in $h_{x}\approx 5$ and $h_{x}\approx 9$. After the second exchange, the 4 Fermi points with smallest wavevectors shift in the same direction when $h_{y}$ is applied, while the other 6 points shift oppositely. Consequently, the top 2 occupied subbands contribute better due to the nonlinear dispersion, and ABSs can balance MBSs better.
S-6 Band topology and exchange condition in multiband Josephson diode
[…]
$\nu=\frac{1}{\pi}\operatorname{Im}\left[\log\left(\prod_{i=1}^{s}\frac{\left|U^{(i)}\right|}{\sqrt{\left|U^{(i)}\right|\left|U^{(i)}\right|^{*}}}\right)\right].$ (S12)
Here, $|\cdot|$ denotes the determinant of the tensor U, and periodic boundary condition should be adopted.
For the ideal 1-orbit model, the criterion for topological phase transition has already been found Liu et al. (2024):
$\mu_{0}^{2}+\Delta^{2}=(h_{x}^{c})^{2}+h_{y}^{2}.$ (S13)
For the 2-orbit case, since $|\mu_{1}-\mu_{2}|\gg\Delta$ in experiment and $h_{y}<\Delta$, when ignoring $\alpha_{y}$, we can find that at $k_{x}=0$, as $h_{x}$ exceeds $h_{x}^{c}\approx\sqrt{\Delta^{2}-h_{y}^{2}}$ and keeps increasing, there will be no band closure. Therefore, no topological phase transition will occur subsequently. Considering $\alpha_{y}$, it only affects the position of $h_{x}^{c}$ to a certain extent, but does not significantly influence the value of the starting point $h_{x}^{c^{\prime}}\approx|\mu_{1}-\mu_{2}|/2$ of the high-efficiency platform.